TERAHERTZ TIME DOMAIN SPECTROSCOPY OF COMPLEX ORGANIC MOLECULES IN ASTRO-PHYSICALLY RELEVANT ICES

\underline{BRETT A. McGUIRE}, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125; SERGIO IOPPOLO, Division of Geological and Planetary Sciences, California Institute of Technology, Pasadena, CA 91125; MARCO A. ALLODI, MATTHEW J. KELLEY, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125; GEOFFREY A. BLAKE, Division of Chemistry and Chemical Engineering and Division of Geological and Planetary Sciences, California Institute of Technology, Pasadena, CA 91125.

We have constructed a new system to study the spectra of astrophysically-relevant ice analogs using THz time-domain spectroscopy from 300 GHz - 7 THz. Here, we present our initial efforts to explore the spectra of pure ices of simple, abundant interstellar species as well as complex organic molecules (COMs) and COM-doped ice mixtures. We will present preliminary spectra of pure H_{2}O, CO_{2}, methanol (CH_{3}OH), and methyl formate (CH_{3}COOH) ices, as well as spectra of these molecules embedded in a variety of other relevant interstellar analogs. Our results are discussed in the context of astronomical observations and the possibility of probing ice compositions in the absence of a background radiation source.